### v3

@@ 4.2 Variable Smoothing Algorithm with Convergence Analysis
    Choose a subsequence $(\bm{y}_{m(l)})_{l=1}^{\infty}\subset\mathcal{Y}$ of $(\bm{y}_{n})_{n=1}^{\infty}$ such that $\lim_{l\to\infty}\left\lVert\nabla(F_{m(l)}\circ\bm{\varphi})(\bm{y}_{m(l)})\right\rVert=0$, where $m:\mathbb{N}\to\mathbb{N}$ is monotonically increasing [^10] From Theorem 4.9 (b), we can construct such a subsequence $(\bm{y}_{m(l)})_{l=1}^{\infty}$ by using, e.g., $m(l)\coloneqq\begin{cases}1&(l=1)\\ \min\{n\in\mathbb{N}\mid n>m(l-1),\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert\leq 2^{-l}\}&(l>1).\end{cases}$ (53) . Then, every cluster point $\bm{y}^{\star}\in\mathcal{Y}$ of $(\bm{y}_{m(l)})_{l=1}^{\infty}$ is a stationary point of Problem 1.2 in the sense of (26).
@@
    where $\chi\in\mathbb{R}_{++}$ is given in Theorem 4.9 (a), and $\lceil\cdot\rceil$ stands for the ceiling function (see also Remark 4.10 below for an interpretation of this result).
@@
The proof for (a) is inspired partially by that for [23, Thm. 4.1].
@@
(b) Let $a_{n}\coloneqq\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert$, and $b_{n}\coloneqq\inf\{a_{m}\mid m\geq n\}$. For every $\underline{k}\in\mathbb{N}$, the inequality (51) with $\overline{k}\to\infty$ implies $b_{\underline{k}}=0$ thanks to the nonsummablility of $(\mu_{n})_{n=1}^{\infty}$ (see the condition (ii) in (41)). This implies $\liminf_{n\to\infty}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert=\liminf_{n\to\infty}a_{n}=\lim_{n\to\infty}b_{n}=0$.
@@
Remark 4.10 (Interpretation of Theorem 4.9 (d)).
For a given $\epsilon>0$, a point $\widehat{\bm{y}}\in\mathcal{Y}$ satisfying $\mu<\epsilon$ and $\left\lVert\nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})(\widehat{\bm{y}})\right\rVert<\epsilon$ with $F^{\langle\mu\rangle}=h+{}^{\mu}g\circ\mathfrak{S}$ is called an $\epsilon$-approximate stationary point [75, 41, 39] of Problem 1.2, where $F^{\langle\mu\rangle}\circ\bm{\varphi}$ has the gradient consistency property for $F\circ\bm{\varphi}$ (see Theorem 4.4 (b) and Remark 4.5). By Theorem 4.9 (d), for a given $\epsilon\in(0,1)$, a generated sequence $(\bm{y}_{n})_{n=1}^{\infty}$ by Algorithm 1 with $\mu_{n}\coloneqq\tau n^{-1/\alpha}\ (\tau\in(0,2\eta)^{-1}],\ \alpha>1)$ achieves an $\epsilon$-approximate stationary point of Problem 1.2 within $\mathcal{O}(\epsilon^{\min\{-2\alpha/(\alpha-1),-\alpha\}})$ iterations, where $\mathcal{O}$ stands for the Landau’s Big O notation. Among $\mathcal{O}(\epsilon^{\min\{-2\alpha/(\alpha-1),-\alpha\}})$ for $\alpha>1$, the parameter $\alpha\coloneqq 3$ achieves the best iteration complexity $\mathcal{O}(\epsilon^{-3})$. The same complexity $\mathcal{O}(\epsilon^{-3})$ is shown for a smoothing method in [41] for the minimization of a nonconvex nonsmooth function $h(\bm{x})+\sum_{i=1}^{N}\psi(\lvert[\bm{x}]_{i}\rvert^{p})\ (\bm{x}\in\mathbb{R}^{N})$, where $h$ is continuously differentiable, and its gradient is Lipschitz continuous; $0<p\leq 1$; and $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is a penalty function such that $\psi$ is continuously differentiable, nondecreasing, and $\psi^{\prime}$ is locally Lipschitz continuous. We note that (i) the parameter $\alpha\coloneqq 3$ in $\mu_{n}\coloneqq\tau n^{-1/\alpha}$ has been utilized in variable smoothing-type algorithms [23, 25, 24] as a reasonable smoothing parameter; and (ii) by Theorem 4.9 (c), an asymptotic convergence property of Algorithm 1 is guaranteed with smoothing parameters $(\mu_{n})_{n=1}^{\infty}$ satisfying (41) other than $\mu_{n}\coloneqq\tau n^{-1/\alpha}$.
@@ 5 Numerical experiments
Both applications are reduced to nonsmooth optimization problems over an embedded submanifold, say the Stiefel manifold ${\rm St}(p,N)\coloneqq\{\bm{U}\in\mathbb{R}^{N\times p}\mid\bm{U}^{\mathsf{T}}\bm{U}=\bm{I}_{p}\}\ (p\leq N)$, of $\mathbb{R}^{N\times p}$ [30, Section 3.3.2]. In order to apply Algorithm 1 to such problems, we employ the generalized inverse Cayley transform, denoted by $\Phi_{\bm{S}}^{-1}$ in [54], as a parametrization $\bm{\varphi}$ of ${\rm St}(p,N)$:
@@
is the set of skew-symmetric matrices, $E_{N,p}(\bm{S})\coloneqq\{\bm{U}\in{\rm St}(p,N)\mid\det(\bm{I}_{p}+\bm{I}_{N\times p}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{U})=0\}$ is called the singular-point set, and $\bm{I}_{N\times p}\in\mathbb{R}^{N\times p}$ is the first $p$ column vectors of the NN-by-NN identity matrix $\bm{I}$. Although the image $\bm{\varphi}_{\bm{S}}^{\rm Cay}(Q_{N,p})={\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ is a proper subset of ${\rm St}(p,N)$, ${\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ is open [^13] There exists an open subset $O\subset\mathbb{R}^{N\times p}$ in the Euclidean space $\mathbb{R}^{N\times p}$ such that ${\rm St}(p,N)\setminus E_{N,p}(\bm{S})=O\cap{\rm St}(p,N)$. and dense [^14] For any $\bm{U}\in{\rm St}(p,N)$, there exists $(\bm{U}_{n})_{n=1}^{\infty}\subset{\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ such that $\lim_{n\to\infty}\bm{U}_{n}=\bm{U}$. in ${\rm St}(p,N)$ [54, Thm. 2.3 (b)], implying thus almost all points in ${\rm St}(p,N)$ can be parametrized in terms of the Euclidean space $Q_{N,p}$ with $\bm{\varphi}_{\bm{S}}^{\rm Cay}$.
Corollary 5.1 below implies that $\bm{\varphi}_{\bm{S}}^{\rm Cay}$ possesses well-suited properties for its application to Algorithm 1 as a parametrization $\bm{\varphi}$. More precisely, with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$, (i) Corollary 5.1 (a) implies the equivalence of the first-order optimality conditions for Problem 1.1 and Problem 1.2; (ii) Corollary 5.1 (b) and Theorem 4.9 imply that Algorithm 1 is guaranteed to approximate a stationary point of Problem 1.2.

### v4

@@ 4.2 Variable Smoothing Algorithm with Convergence Analysis
    Choose a subsequence $(\bm{y}_{m(l)})_{l=1}^{\infty}\subset\mathcal{Y}$ of $(\bm{y}_{n})_{n=1}^{\infty}$ such that $\lim_{l\to\infty}\left\lVert\nabla(F_{m(l)}\circ\bm{\varphi})(\bm{y}_{m(l)})\right\rVert=0$, where $m:\mathbb{N}\to\mathbb{N}$ is monotonically increasing [^11] From Theorem 4.9 (b), we can construct such a subsequence $(\bm{y}_{m(l)})_{l=1}^{\infty}$ by using, e.g., $m(l)\coloneqq\begin{cases}1&(l=1)\\ \min\{n\in\mathbb{N}\mid n>m(l-1),\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert\leq 2^{-l}\}&(l>1).\end{cases}$ (54) . Then, every cluster point $\bm{y}^{\star}\in\mathcal{Y}$ of $(\bm{y}_{m(l)})_{l=1}^{\infty}$ is a stationary point of Problem 1.4. Moreover, if $N_{C}(\bm{\varphi}(\bm{y}^{\star}))\supset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$ holds, then $\bm{\varphi}(\bm{y}^{\star})\in C$ is a stationary point of Problem 1.1 by Theorem 3.2 (b) (see also Corollary 3.4 and Remark 3.6).
@@
    where $\chi\in\mathbb{R}_{++}$ is given in Theorem 4.9 (a), $\lceil\cdot\rceil$ stands for the ceiling function, and such $\bm{y}_{n}$ is an $\epsilon$-approximate stationary point defined in Definition 4.10 (b).
@@
The proof for (a) is inspired partially by that for [13, Thm. 4.1].
@@
(b) Let $a_{n}\coloneqq\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert$, and $b_{n}\coloneqq\inf\{a_{m}\mid m\geq n\}$. For every $\underline{k}\in\mathbb{N}$, the inequality (52) with $\overline{k}\to\infty$ implies $b_{\underline{k}}=0$ thanks to the nonsummability of $(\mu_{n})_{n=1}^{\infty}$ (see the condition (ii) in (42)). This implies $\liminf_{n\to\infty}\left\lVert\nabla(F_{n}\circ\bm{\varphi})(\bm{y}_{n})\right\rVert=\liminf_{n\to\infty}a_{n}=\lim_{n\to\infty}b_{n}=0$.
@@
To derive a first-order oracle complexity of Algorithm 1, we introduce the definitions of the first-order oracle and an $\epsilon$-approximate stationary point, e.g., [37, 12, 88], for smoothing-type algorithms.
Definition 4.10.
Consider Problem 1.4 with $F^{\langle\mu\rangle}=h+{}^{\mu}g\circ\mathfrak{S}$.
  1. (a)
    For given $\bm{x},\bm{u}\in\mathcal{X}$, $\bm{y}\in\mathcal{Y}$, and $\bm{z},\bm{v}\in\mathcal{Z}$, and $\mu\in(0,\eta^{-1})$, the first-order oracle returns $h(\bm{x})$, $\nabla h(\bm{x})$, $\mathfrak{S}(\bm{x})$, $(\mathrm{D}\mathfrak{S}(\bm{x}))^{*}[\bm{v}]$, $\bm{\varphi}(\bm{y})$ $(\mathrm{D}\bm{\varphi}(\bm{y}))^{*}[\bm{u}]$, $g(\bm{z})$, and $\mathrm{prox}_{\mu g}(\bm{z})$ (Note: one evaluation of $F^{\langle\widebar{\mu}\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})$ and $\nabla(F^{\langle\widebar{\mu}\rangle}\circ\bm{\varphi})(\widebar{\bm{y}})$ for $\widebar{\bm{y}}\in\mathcal{Y}$ and $\widebar{\mu}\in(0,\eta^{-1})$ can be implemented with at most five oracle calls [^13] By noting that * denotes an arbitrary value, consider the following five oracle calls with (i) $(\bm{x},\bm{u},\bm{y},\bm{z},\bm{v},\mu)=(*,*,\widebar{\bm{y}},*,*,*)$ to obtain $\widebar{\bm{x}}\coloneqq\bm{\varphi}(\widebar{\bm{y}})\in\mathcal{X}$; (ii) $(\bm{x},\bm{u},\bm{y},\bm{z},\bm{v},\mu)=(\widebar{\bm{x}},*,*,*,*,*)$ to obtain $\widebar{\xi}_{\rm cost}\coloneqq h(\widebar{\bm{x}})\in\mathbb{R}$, $\widebar{\bm{\xi}}_{\rm grad}\coloneqq\nabla h(\widebar{\bm{x}})\in\mathcal{X}$, $\widebar{\bm{z}}\coloneqq\mathfrak{S}(\widebar{\bm{x}})\in\mathcal{Z}$; (iii) $(\bm{x},\bm{u},\bm{y},\bm{z},\bm{v},\mu)=(*,*,*,\widebar{\bm{z}},*,\widebar{\mu})$ to obtain $\widebar{\bm{p}}\coloneqq\mathrm{prox}_{\widebar{\mu}g}(\widebar{\bm{z}})\in\mathcal{Z}$; (iv) $(\bm{x},\bm{u},\bm{y},\bm{z},\bm{v},\mu)=(\widebar{\bm{x}},*,*,\widebar{\bm{p}},(\widebar{\bm{z}}-\widebar{\bm{p}})/\widebar{\mu},*)$ to obtain $\widebar{\zeta}_{\rm cost}\coloneqq g(\widebar{\bm{p}})\in\mathbb{R}$, $\widebar{\bm{\zeta}}_{\rm grad}\coloneqq\left(\mathrm{D}\mathfrak{S}(\widebar{\bm{x}})\right)^{*}[(\widebar{\bm{z}}-\widebar{\bm{p}})/\widebar{\mu}]\in\mathcal{X}$; (v) $(\bm{x},\bm{u},\bm{y},\bm{z},\bm{v},\mu)=(*,\widebar{\bm{\xi}}_{\rm grad}+\widebar{\bm{\zeta}}_{\rm grad},\widebar{\bm{y}},*,*,*)$ to obtain $\nabla(F^{\langle\widebar{\mu}\rangle}\circ\bm{\varphi})(\widebar{\bm{y}})=(\mathrm{D}\bm{\varphi}(\widebar{\bm{y}}))^{*}[\widebar{\bm{\xi}}_{\rm grad}+\widebar{\bm{\zeta}}_{\rm grad}]\in\mathcal{Y}$, where we can compute $F^{\langle\widebar{\mu}\rangle}\circ\bm{\varphi}(\widebar{\bm{y}})=\widebar{\xi}_{\rm cost}+\widebar{\zeta}_{\rm cost}+\frac{1}{2\widebar{\mu}}\left\lVert\widebar{\bm{p}}-\widebar{\bm{z}}\right\rVert^{2}\in\mathbb{R}$ (Compare with the expression (28)). ).
  2. (b)
    For a given $\epsilon>0$, a point $\widehat{\bm{y}}\in\mathcal{Y}$ is said to be an $\epsilon$-approximate stationary point of Problem 1.4 if $\left\lVert\nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})(\widehat{\bm{y}})\right\rVert\leq\epsilon$ holds for some $\mu\in(0,\epsilon]$
Compared with the oracle in [83] proposed for Problem 1.1 with $g$ convex, our first-order oracle returns additionally the values $\bm{\varphi}(\bm{y})$ and its derivative $(\mathrm{D}\bm{\varphi}(\bm{y}))^{*}[\bm{u}]$ for $\bm{y}\in\mathcal{Y}$ and $\bm{u}\in\mathcal{X}$, where an evaluation of $\bm{\varphi}$ corresponds to that of a retraction used in [83] for mapping a tangent vector to a point in the manifold $C$ (see just after (90) for a retraction). Moreover, our oracle is essentially the same one considered in [13] for Problem 1.4 in a special case where $\bm{\varphi}=\mathrm{Id}$ and $\mathfrak{S}$ is linear.
Our approximate stationarity notion for Problem 1.4 in Definition 4.10 (b) is not directly comparable to the notions used in, [66, 10, 83] and [25, 80], for Problem 1.1. Instead, our approximate stationarity reduces to another notion of approximate stationarity used in [13] for Problem 1.4 with $\bm{\varphi}=\mathrm{Id}$ and $\mathfrak{S}$ linear (see Remark 4.11).
Remark 4.11 (Relation between approximate stationarities in Definition 4.10 (b) and in [13]).
For Problem 1.4, another type of an $\epsilon$-approximate stationary point can be defined as a point $\widehat{\bm{y}}\in\mathcal{Y}$ satisfying
$\exists\bm{z}\in\mathcal{Z}\ \mathrm{s.t.}\ d\left(-\nabla(h\circ\bm{\varphi})(\widehat{\bm{y}}),(\mathrm{D}(\mathfrak{S}\circ\bm{\varphi})(\widehat{\bm{y}}))^{*}\left[\partial_{\rm L}g(\bm{z})\right]\right)\leq\epsilon\ \mathrm{and}\ \left\lVert\bm{z}-\mathfrak{S}\circ\bm{\varphi}(\widehat{\bm{y}})\right\rVert\leq\epsilon.$ (73)
In a special case where $\bm{\varphi}\coloneqq\mathrm{Id}$ and $\mathfrak{S}$ is linear, a variable smoothing algorithm in [13] can find an $\epsilon$-approximate stationary point in the sense of (73) within $\mathcal{O}(\epsilon^{-3})$ oracle calls. In fact, our approximate stationary point in Definition 4.10 (b) implies an approximate stationary point in the sense of (73) as follows. Let $\widehat{\bm{y}}\in\mathcal{Y}$ be an $\epsilon$-approximate stationary point in Definition 4.10 (b), i.e., $\left\lVert\nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})(\widehat{\bm{y}})\right\rVert\leq\epsilon$ with some $\mu\in(0,\epsilon]$. Then, we have
[…]
$=$ $\left\lVert\nabla(h\circ\bm{\varphi})(\widehat{\bm{y}})+(\mathrm{D}(\mathfrak{S}\circ\bm{\varphi})(\widehat{\bm{y}}))^{*}[\nabla{}^{\mu}g(\mathfrak{S}\circ\bm{\varphi}(\widehat{\bm{y}}))]\right\rVert=\left\lVert\nabla(F^{\langle\mu\rangle}\circ\bm{\varphi})(\widehat{\bm{y}})\right\rVert\leq\epsilon.$ (76)
Moreover, with the Lipschitz constant $L_{g}$ of $g$, $\left\lVert\mathrm{prox}_{\mu g}\left(\mathfrak{S}\circ\bm{\varphi}(\widehat{\bm{y}})\right)-\mathfrak{S}\circ\bm{\varphi}(\widehat{\bm{y}})\right\rVert\leq L_{g}\mu$ holds from $\left\lVert\mathrm{prox}_{\mu g}(\widebar{\bm{z}})-\widebar{\bm{z}}\right\rVert\leq L_{g}\mu\ (\widebar{\bm{z}}\in\mathcal{Z})$ (see Fact 2.7 (b) and (22)). Hence, $\widehat{\bm{y}}$ is a $(\max\{1,L_{g}\}\epsilon)$-approximate stationary point in the sense of (73).
In Corollary 4.12, we derive the first-order complexity, i.e., the number of oracle calls, $\mathcal{O}(\epsilon^{-3})$ of Algorithm 1 for finding an $\epsilon$-approximate stationary point of Problem 1.4. We note that, under a special case where $\bm{\varphi}=\mathrm{Id}$ and $\mathfrak{S}$ is linear, Corollary 4.12 recovers the oracle complexity $\mathcal{O}(\epsilon^{-3})$ of the existing variable smoothing algorithm [13] (see Remark 4.11 and Table 1 in Section 1), whereas the proposed algorithm admits the nonlinearity of $\mathfrak{S}$.
Corollary 4.12 (First-order oracle complexity of Algorithm 1).
Consider Problem 1.4. Choose arbitrarily $\bm{y}_{1}\in\mathcal{Y}$, $c\in(0,1)$, and $\mu_{n}\coloneqq\tau n^{-1/3}\ (n\in\mathbb{N})$ with $\tau\in(0,(2\eta)^{-1}]$. Then, under Assumption 4.7, for any given $\epsilon>0$, Algorithm 1 can find an $\epsilon$-approximate stationary point of Problem 1.4 (i) with at most $\mathcal{O}(\epsilon^{-3})$ first-order oracle calls if $\gamma_{n}\coloneqq 2(1-c)L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}\ (n\in\mathbb{N})$ (see (48) and Example 4.8 (a)), and (ii) with at most $\mathcal{O}(\epsilon^{-3}\log(\epsilon^{-1}))$ first-order oracle calls if $\gamma_{n}$ is given by Algorithm 2 with $\rho\in(0,1)$ and $\gamma_{\rm init}\in\mathbb{R}_{++}$ (see Example 4.8 (b)).
Proof.
Recall that Algorithm 1 can find an $\epsilon$-approximate stationary point within $\mathcal{O}(\epsilon^{-3})$ iterations by Theorem 4.9 (d) with $\alpha\coloneqq 3$. Hence, for the case (i) $\gamma_{n}\coloneqq 2(1-c)L_{\nabla(F_{n}\circ\bm{\varphi})}^{-1}\ (n\in\mathbb{N})$, Algorithm 1 can find an $\epsilon$-approximate stationary point with at most $\mathcal{O}(\epsilon^{-3})$ first-order oracle calls because $F_{n}\circ\bm{\varphi}$ and $\nabla(F_{n}\circ\bm{\varphi})$ are evaluated only once at each iteration respectively (see also Example 4.8 (a)).
For the case (ii), at $n$th iteration of Algorithm 1, Algorithm 2 for estimating stepsize $\gamma_{n}$ terminates within $\max\left\{1,\left\lceil\log_{\rho}\left(\frac{2\mu_{n}(1-c)}{\gamma_{\rm init}(\mu_{n}\varpi_{1}+\varpi_{2})}\right)\right\rceil\right\}$ updates (see Example 4.8 (b)). Since $F_{n}\circ\bm{\varphi}$ is computed once at each update of $\gamma_{n}$ in Algorithm 2, $F_{n}\circ\bm{\varphi}$ is evaluated for $M_{n}\coloneqq\max\left\{1,\left\lceil\log_{\rho}\left(\frac{2\mu_{n}(1-c)}{\gamma_{\rm init}(\mu_{n}\varpi_{1}+\varpi_{2})}\right)\right\rceil\right\}+1$ times at most in Algorithm 2 for estimating $\gamma_{n}$ at $n$th iteration of Algorithm 1. Hence, the first-order oracle complexity can be obtained by evaluating an upper bound of $\sum_{n=1}^{\mathcal{O}(\epsilon^{-3})}M_{n}$. Since $\log_{\rho}(\cdot)$ with $\rho\in(0,1)$ is monotonically decreasing and $\mu_{n}=\tau n^{-1/3}\leq\tau\ (n\in\mathbb{N})$, we have $\log_{\rho}(\frac{2\mu_{n}(1-c)}{\gamma_{\rm init}(\mu_{n}\varpi_{1}+\varpi_{2})})\leq\log_{\rho}(\frac{2\mu_{n}(1-c)}{\gamma_{\rm init}(\tau\varpi_{1}+\varpi_{2})})=\log_{\rho}(\frac{2\tau(1-c)n^{-1/3}}{\gamma_{\rm init}(\tau\varpi_{1}+\varpi_{2})})=-\frac{1}{3}\log_{\rho}(n)+\Delta=-\frac{\log(n)}{3\log(\rho)}+\Delta$ with some $\Delta\in\mathbb{R}$. From $\lceil t\rceil\leq t+1\leq\lvert t\rvert+1\ (t\in\mathbb{R})$, we get $M_{n}\leq\lvert\log_{\rho}(\frac{2\mu_{n}(1-c)}{\gamma_{\rm init}(\mu_{n}\varpi_{1}+\varpi_{2})})\rvert+1+1\leq\frac{\log(n)}{3\lvert\log(\rho)\rvert}+\lvert\Delta\rvert+2$. From $\sum_{n=1}^{k}\log(n)\leq\sum_{n=1}^{k}\log(k)=k\log(k)\ (k\in\mathbb{N})$, we get $\sum_{n=1}^{k}M_{n}\leq\frac{k\log(k)}{3\lvert\log(\rho)\rvert}+k(\lvert\Delta\rvert+2)$. By substituting $k=\mathcal{O}(\epsilon^{-3})$, we get the first-order oracle complexity $\mathcal{O}(\epsilon^{-3}\log(\epsilon^{-3}))=\mathcal{O}(\epsilon^{-3}\log(\epsilon^{-1}))$. ∎
@@ 5 Numerical experiments
Both applications are reduced to nonsmooth optimization problems over an embedded submanifold, say the Stiefel manifold ${\rm St}(p,N)\coloneqq\{\bm{U}\in\mathbb{R}^{N\times p}\mid\bm{U}^{\mathsf{T}}\bm{U}=\bm{I}_{p}\}\ (p\leq N)$, of $\mathbb{R}^{N\times p}$ [3, Section 3.3.2]. In order to apply Algorithm 1 to such problems, we employ the generalized inverse Cayley transform, denoted by $\Phi_{\bm{S}}^{-1}$ in [48], as a parametrization $\bm{\varphi}$ of ${\rm St}(p,N)$:
@@
is the set of skew-symmetric matrices, $E_{N,p}(\bm{S})\coloneqq\{\bm{U}\in{\rm St}(p,N)\mid\det(\bm{I}_{p}+\bm{I}_{N\times p}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{U})=0\}$ is called the singular-point set, and $\bm{I}_{N\times p}\in\mathbb{R}^{N\times p}$ is the first $p$ column vectors of the NN-by-NN identity matrix $\bm{I}$. Although the image $\bm{\varphi}_{\bm{S}}^{\rm Cay}(Q_{N,p})={\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ is a proper subset of ${\rm St}(p,N)$, ${\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ is open [^15] There exists an open subset $O\subset\mathbb{R}^{N\times p}$ in the Euclidean space $\mathbb{R}^{N\times p}$ such that ${\rm St}(p,N)\setminus E_{N,p}(\bm{S})=O\cap{\rm St}(p,N)$. and dense [^16] For any $\bm{U}\in{\rm St}(p,N)$, there exists $(\bm{U}_{n})_{n=1}^{\infty}\subset{\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ such that $\lim_{n\to\infty}\bm{U}_{n}=\bm{U}$. in ${\rm St}(p,N)$ [48, Thm. 2.3 (b)], implying thus almost all points in ${\rm St}(p,N)$ can be parameterized in terms of the Euclidean space $Q_{N,p}$ with $\bm{\varphi}_{\bm{S}}^{\rm Cay}$.
Corollary 5.1 below implies that $\bm{\varphi}_{\bm{S}}^{\rm Cay}$ possesses well-suited properties for its application to Algorithm 1 as a parametrization $\bm{\varphi}$. More precisely, with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$, (i) Corollary 5.1 (a) implies the equivalence of the first-order optimality conditions for Problem 1.1 and Problem 1.4; (ii) Corollary 5.1 (b) and Theorem 4.9 imply that Algorithm 1 is guaranteed to approximate a stationary point of Problem 1.4.